# LivePyxel: Accelerating image annotations with a Python-integrated webcam live streaming Uriel Garcilazo-Cruz,a,b,†, Joseph O. Okemea,b, and Rodrigo A. Vargas-Hernándeza,b,c,†

### v1

@@ 3.1 Water Tank: Image Segmentation Task
The presence of certain microorganisms in the environment often serves as an indirect indicator of ecosystem health 31. These organisms, commonly referred to as indicator species (IS), are highly sensitive to chemical and physical changes in their environment, and their use has become widespread in environmental assessment studies 32. Some datasets and data management digital infrastructures have been published in recent years 33, 34, 35, highlighting the growing importance of automation in the identification and monitoring of these species communities. However, the ephemeral nature of these communities poses significant challenges in collecting and processing annotations for vision-model tasks at a rate comparable to the speed at which such communities change. To demonstrate the capabilities of LivePyxel, we curated a high-quality dataset capturing the biodiversity of water samples collected from an urban park in Toronto, Canada, in April 2025. We acquired 1,250 image-mask pairs from a water tank, belonging to eight categories of microorganisms (see Supplementary Material for more information).
@@
Fig.  6: The F1 scores achieved by a U-Net model, highlighting the performance across the eight different microorganisms and the background. The U-Net was initialized with the VGG-19 weights. For more details regarding the training, see the main text.
Fig.  7: Example predictions from the trained U-Net model on the Water Tank (EM) dataset. Each row shows the original microscopy image (left), the corresponding ground truth mask (middle), and the model’s predicted segmentation (right). The legend on the right indicates the color coding for each class.
@@ 3.2 Snail Shells: Data engineering with binary masks
Creating training datasets for semantic segmentation is notoriously labor-intensive. Precise pixel-level annotations are required, and this process becomes especially difficult for images containing many small and round objects, where annotators must painstakingly outline each instance. For example, producing 1,000 segmentation masks for a benchmark like MS COCO took about 22 37 hours, highlighting the high cost of manual labeling. The performance of most vision models improves significantly with the size of the training dataset. This creates a strong incentive to explore automated and semi-automated annotation methods and advanced data augmentation to expand datasets without proportional human effort 37.
In certain scenarios, a dataset can be constructed under highly controlled lighting and background conditions to semi-automate the annotation process. For example, using a uniform background (e.g., a solid white or green backdrop) and exploiting simple threshold-based image operations to obtain binary masks for the foreground objects. In a controlled setup where the background is a known uniform color (such as white), images can be converted to grayscale (or other color space) and apply a threshold that classifies each pixel as background (0) or object ($1$) based on its intensity or hue 38. This effectively binarizes the image, separating the objects from the backdrop. For example, researchers have captured objects on a plain white background under consistent lighting, which reduces the complexity of background removal and allows easy discrimination of the object by simple intensity thresholding when dealing with organic shapes 39. In fields like medical imaging, placing surgical instruments in front of a green screen has been used to automatically extract tool masks – the monochromatic green background makes it easy to isolate the instrument by hue thresholding 38. In such controlled conditions, the threshold can be tuned so that any pixel brighter (or darker) than a set value is labeled as background, while the rest are labeled as foreground (object). This semi-automated mask generation dramatically speeds up dataset creation, since the bulk of the annotation is handled by an algorithm (with minimal human correction for any errors).
@@
Fig.  8: Pipeline for Engineered Semantic Segmentation Dataset Creation. (Panel A) The binary mask tool from LivePyxel captures an image/mask annotation of shells belonging to the same class. (Panel B) Masks are overlapped on top of images to retrieve transparencies. (Panel C) Transparencies are randomly stacked together, forming a new composite image on top of a background. (Panel D) We quickly engineered a dataset consisting of 10,000 image/mask pairs.
@@
We implemented a randomized compositing strategy. For each of the 10,000 engineered samples, a base background image was selected and the class folders were shuffled to introduce variation. A random number of classes were sampled, and for each selected class, a transparent image was randomly chosen and placed on the background (Fig. 8- C). The same transformation, such as flipping or rotation, was applied to both the image cutout and its corresponding mask. This ensured that pixel-level alignment was preserved within an image/mask pair. Multiple instances from different classes were overlaid in succession, resulting in composite scenes that mimic realistic configurations with precise pixel-accurate masks for each object (Fig. 8- D).
This pipeline produced a total of 10,000 synthetic image/mask pairs, significantly enriching the dataset and introducing diverse combinations of object instances, orientations, and overlaps. These engineered samples were subsequently used to train the same U-Net segmentation model used for Section 3.1, keeping the image size at 512. The model was trained for 24 hours on 4 NVIDIA GPUs A100. In contrast with the water tank dataset (Section 3.1), the F1 scores for the snail shells dataset demonstrated consistently high performance across all classes, with minimal variation between training and validation sets and a very high F1 score across all categories (Fig. 9). The model also exhibited a rapid decline in the loss function, accompanied by a steep rise in F1 scores within the first three epochs (Fig. S.5 and Fig. S.6 in Supplemental Material). Fig. 10 showcases the predicted masks by the trained U-Net for some of the images in the dataset.
Fig.  9: F1 scores achieved by a U-Net model with VGG-19 backbone during training and validation. The plot highlights performance across 5 different classes.
Fig.  10: Example predictions from the trained U-Net model on the snail shells dataset. Each set of three columns shows the original image (left), the corresponding ground truth mask (middle), and the model’s predicted segmentation (right). The legend below indicates the color coding for each class.
@@ 4 Summary
We present LivePyxel, an open-source pixel-level annotation library designed for both real-time image labeling and pre-saved image datasets. Built on OpenCV, LivePyxel efficiently handles image processing tasks and introduces Bézier splines as a key feature, enabling smoother and more precise mask boundaries. To demonstrate LivePyxel’s utility, we provide examples where the labeled images, generated with LivePyxel, are used for training a vision model, completing the annotation-to-inference loop. The datasets for both presented examples were captured using a GoPro camera mounted on a microscope, highlighting LivePyxel’s compatibility with diverse imaging setups. Finally, LivePyxel allows the use of tablet-type devices, allowing the annotation process to be more ergonomic.
@@
LivePyxel and all trained vision models and data presented in this paper are freely available at https://github.com/UGarCil/LivePyxel.

### v2

@@ 3.1 Water Tank: Image Segmentation Task
The presence of certain microorganisms in the environment often serves as an indirect indicator of ecosystem health cairns1993proposed. These organisms, commonly referred to as indicator species (IS), are highly sensitive to chemical and physical changes in their environment, and their use has become widespread in environmental assessment studies siddig2016ecologists. Some datasets and data management digital infrastructures have been published in recent years li2021emds, li2020developing, goldberg2005open, highlighting the growing importance of automation in the identification and monitoring of these species communities. However, the ephemeral nature of these communities poses significant challenges in collecting and processing annotations for vision-model tasks at a rate comparable to the speed at which such communities change. To demonstrate the capabilities of LivePyxel, we curated a high-quality dataset capturing the biodiversity of water samples collected from an urban park in Toronto, Canada, in April 2025. We acquired 1,250 image-mask pairs from a water tank, belonging to eight categories of microorganisms (see Supplementary Material for more information).
@@
Fig. 5: The F1 scores achieved by a U-Net model, highlighting the performance across the eight different microorganisms and the background. The U-Net was initialized with the VGG-19 weights. For more details regarding the training, see the main text.
Fig. 6: Example predictions from the trained U-Net model on the Water Tank (EM) dataset. Each row shows the original microscopy image (left), the corresponding ground truth mask (middle), and the model’s predicted segmentation (right). The legend on the right indicates the color coding for each class.
@@ 3.2 Snail Shells: Data engineering with binary masks
Creating training datasets for semantic segmentation is notoriously labor-intensive. Precise pixel-level annotations are required, and this process becomes especially difficult for images containing many small and round objects, where annotators must painstakingly outline each instance. For example, producing 1,000 segmentation masks for a benchmark like MS COCO took about 22 ghiasi2021simple hours, highlighting the high cost of manual labeling. The performance of most vision models improves significantly with the size of the training dataset. This creates a strong incentive to explore automated and semi-automated annotation methods and advanced data augmentation to expand datasets without proportional human effort ghiasi2021simple.
In certain scenarios, a dataset can be constructed under highly controlled lighting and background conditions to semi-automate the annotation process. For example, using a uniform background (e.g., a solid white or green backdrop) and exploiting simple threshold-based image operations to obtain binary masks for the foreground objects. In a controlled setup where the background is a known uniform color (such as white), images can be converted to grayscale (or other color space) and apply a threshold that classifies each pixel as background (0) or object ($1$) based on its intensity or hue garcia2021image. This effectively binarizes the image, separating the objects from the backdrop. For example, researchers have captured objects on a plain white background under consistent lighting, which reduces the complexity of background removal and allows easy discrimination of the object by simple intensity thresholding when dealing with organic shapes mohana2015automatic. In fields like medical imaging, placing surgical instruments in front of a green screen has been used to automatically extract tool masks – the monochromatic green background makes it easy to isolate the instrument by hue thresholding garcia2021image. In such controlled conditions, the threshold can be tuned so that any pixel brighter (or darker) than a set value is labeled as background, while the rest are labeled as foreground (object). This semi-automated mask generation dramatically speeds up dataset creation, since the bulk of the annotation is handled by an algorithm (with minimal human correction for any errors).
@@
Fig. 7: Pipeline for Engineered Semantic Segmentation Dataset Creation. (Panel A) The binary mask tool from LivePyxel captures an image/mask annotation of shells belonging to the same class. (Panel B) Masks are overlapped on top of images to retrieve transparencies. (Panel C) Transparencies are randomly stacked together, forming a new composite image on top of a background. (Panel D) We quickly engineered a dataset consisting of 10,000 image/mask pairs.
@@
We implemented a randomized compositing strategy. For each of the 10,000 engineered samples, a base background image was selected, and the class folders were shuffled to introduce variation. A random number of classes were sampled, and for each selected class, a transparent image was randomly chosen and placed on the background (Fig. 7- C). The same transformation, such as flipping or rotation, was applied to both the image cutout and its corresponding mask. This ensured that pixel-level alignment was preserved within an image/mask pair. Multiple instances from different classes were overlaid in succession, resulting in composite scenes that mimic realistic configurations with precise pixel-accurate masks for each object (Fig. 7-D).
This pipeline produced a total of 10,000 synthetic image/mask pairs, significantly enriching the dataset and introducing diverse combinations of object instances, orientations, and overlaps. These engineered samples were subsequently used to train the same U-Net segmentation model used for Section 3.1, keeping the image size at 512. The model was trained for 24 hours on 4 NVIDIA A100 GPUs. In contrast with the water tank dataset (Section 3.1), the F1 scores for the snail shells dataset demonstrated consistently high performance across all classes, with minimal variation between training and validation sets and a very high F1 score across all categories (Fig. 8). The model also exhibited a rapid decline in the loss function, accompanied by a steep rise in F1 scores within the first three epochs (Fig. S.5 and Fig. S.6 in Supplemental Material). Fig. 9 showcases the predicted masks by the trained U-Net for some of the images in the dataset.
Fig. 8: F1 scores achieved by a U-Net model with VGG-19 backbone during training and validation. The plot highlights performance across 5 different classes.
Fig. 9: Example predictions from the trained U-Net model on the snail shells dataset. Each set of three columns shows the original image (left), the corresponding ground truth mask (middle), and the model’s predicted segmentation (right). The legend below indicates the color coding for each class.
@@
3.3 Comparison between polygon and Bézier-splines
We evaluated the performance of LivePyxel against several widely used annotation tools. To ensure a fair comparison, a reference image composed of well-defined polygons with varying degrees of curvature complexity was generated (panel labeled "Original" in Fig. 10).
Our results show that the accuracy achieved with LivePyxel is comparable to, or slightly higher than, other annotation software packages, while maintaining similar annotation times; see Section 5 in the SM. As highlighted in Section 3.2, a key advantage of LivePyxel is its capability to perform Boolean operations directly within the mask. This feature, uncommon among existing tools, enables rapid and flexible labeling of complex regions, including those with internal holes like the central object in Fig. 10.
In terms of performance, LivePyxel exhibited a balanced trade-off between false positives (5.7%) and false negatives (0.5%), comparable to other tools (Fig. 10). CVAT achieved the lowest overall error (2.5% false positives, 1.3% false negatives) through AI-assisted segmentation, but at the cost of a more complex setup, internet dependency, and non-local data handling. VIA offered the simplest installation (a standalone HTML file), whereas LabelMe required manual dependency management on some systems. COCO Annotator had the most challenging setup, involving Docker and SQL-based database configuration. Among these, only LivePyxel integrates Bézier-spline support, providing smoother boundary representation than polygon-based tools such as VIA, LabelMe, and COCO Annotator.
Fig. 10: Comparison of annotation accuracy across different annotation tools using a reference image (Original) with various levels of curvature complexity. Blue regions represent false positives (areas incorrectly included), while red regions denote false negatives (areas missed). Percentages in each panel indicate the total proportion of false positives (blue) and false negatives (red) for each tool relative to the total number of pixels in the image.
@@ 4 Summary
We present LivePyxel, an open-source pixel-level annotation library designed for both on-site image labeling and pre-saved image datasets. Built on OpenCV, LivePyxel efficiently handles image processing tasks and introduces Bézier splines as a key feature, enabling smoother and more precise mask boundaries. To demonstrate LivePyxel’s utility, we provide examples where the labeled images, generated with LivePyxel, are used for training a vision model, completing the annotation-to-inference loop. The datasets for both presented examples were captured using a GoPro camera mounted on a microscope, highlighting LivePyxel’s compatibility with diverse imaging setups. Finally, LivePyxel supports the use of tablet-type devices, making the annotation process more ergonomic. In the near future, we aim to implement two additional features, i) zoom integration, allowing users to work on finer details within LivePyxel’s central canvas (Fig. 2-E), and ii) inspired by CVAT’s AI integration tools, adding a tool capable of loading AI vision models like SAM to help accelerate the annotation process. Even in its present form, LivePyxel represents a significant innovation for labeling images, addressing an unmet need for smoother editing tools that could accelerate data generation.
@@
LivePyxel and all trained vision models and data presented in this paper are freely available at https://github.com/UGarCil/LivePyxel and can be installed through PyPI.